# Denoising Vision Transformers

### v1

@@ Abstract
We delve into a nuanced but significant challenge inherent to Vision Transformers (ViTs): feature maps of these models exhibit grid-like artifacts (“Original” in Figure 1), which detrimentally hurt the performance of ViTs in downstream tasks. Our investigations trace this fundamental issue down to the positional embeddings at the input stage. To address this, we propose a novel noise model, which is universally applicable to all ViTs. Specifically, the noise model dissects ViT outputs into three components: a semantics term free from noise artifacts and two artifact-related terms that are conditioned on pixel locations. Such a decomposition is achieved by enforcing cross-view feature consistency with neural fields in a per-image basis. This per-image optimization process extracts artifact-free features from raw ViT outputs, providing clean features for offline applications. Expanding the scope of our solution to support online functionality, we introduce a learnable denoiser to predict artifact-free features directly from unprocessed ViT outputs, which shows remarkable generalization capabilities to novel data without the need for per-image optimization. Our two-stage approach, termed Denoising Vision Transformers (DVT), does not require re-training existing pre-trained ViTs and is immediately applicable to any Transformer-based architecture. We evaluate our method on a variety of representative ViTs (DINO, MAE, DeiT-III, EVA02, CLIP, DINOv2, DINOv2-reg). Extensive evaluations demonstrate that our DVT consistently and significantly improves existing state-of-the-art general-purpose models in semantic and geometric tasks across multiple datasets (e.g., +3.84 mIoU). We hope our study will encourage a re-evaluation of ViT design, especially regarding the naive use of positional embeddings.
Figure 1: Denoising Vision Transformers (DVT) removes the noisy artifacts in visual features present in almost all Vision Transformers (ViTs). We use a representative set of ViTs as examples, including supervised (e.g. DeiT-III [32], Auto-aug ViT [10, 29]), reconstruction (e.g., EVA-02 [13]), self-distillation (e.g., DINOv2 [22], DINOv2-reg [8]), and multi-modal (e.g., CLIP [26]) algorithms. Top: Each image triplet showcases an input image, its corresponding raw feature visualization, and the cleaned feature map denoised by DVT. Bottom: These triplets display, in order, a feature map, a K-Means cluster map, and a similarity map of the central patch (red dotted) with other patches in the image. Observe how the artifacts negatively impact clustering accuracy and similarity correspondences and how our DVT effectively addresses these issues. The feature colors in the visualizations are produced using principle component analysis (PCA). Best viewed in color.
@@ 1 Introduction
In recent years, Transformers [34] have emerged as the universal architecture for modern foundation models across many modalities, from language to audio [19, 36], text [27, 6, 24, 1], and images [10, 2]. Vision Transformers (ViTs) [10] are now the new de-facto standard in vision-related tasks. These models not only achieve state-of-the-arts under multiple benchmarks but also exhibit intriguing behaviors and capabilities across various tasks [4, 15, 26, 22].
@@
Despite these significant strides made by ViTs, our work reveals a crucial yet often overlooked challenge: the presence of persistent noise artifacts in ViT outputs, observable across various training algorithms [10, 22, 32, 26, 13, 15, 4] (illustrated in Figure 1). These artifacts, beyond being visually annoying, hinder feature interpretability and disrupt semantic coherence. For example, the bottom row of Figure 1 demonstrates that applying clustering algorithms directly on the raw ViT outputs results in noisy clusters. This issue, prevalent across numerous existing pre-trained ViTs, hinders model performance in downstream tasks, underscoring the need for a complete study to mitigate these artifacts. To that end, this paper aims to answer a crucial research question: Is it feasible to effectively denoise these artifacts in pre-trained ViTs, ideally without model re-training?
@@
To answer this, we first investigate the origins of these artifacts. We posit that positional embeddings, a fundamental component of ViT architecture, significantly contribute to this phenomenon. Our initial analysis substantially supports this hypothesis: First, when a zero-tensor (i.e., no content) is fed into a pre-trained DINOv2 model [22], the resulting output is predominantly characterized by similar noise patterns (Figure 2-(a, 2)). Second, we observe a notable absence of such artifacts in the outputs of a DINOv2 model trained without positional embeddings, which contrasts sharply with the standard model outputs (Figure 2-(a, 1) v.s. (a, 3)). Finally, despite the significant differences in the context of various input frames, the artifacts maintain a consistent relative position in the images (Figure 2-(c), middle row).
Figure 2: Impact of positional embeddings in ViTs. (a) Comparison between DINOv2 ViTs [22] trained with and without positional embeddings ((“ViT” v.s. “$*{}^{*}$”), showcasing feature maps for: (1) a standard ViT process, (2) ViT using only positional embeddings (PE) as input, emphasizing the emergence of artifacts, and (3) a PE-free $*{}^{*}$ process, displaying a clear absence of these artifacts. In the figure, “Patch”: patch embedding, “PE”: position embedding. (b) Illustration of how ViT retains and propagates the positional embeddings. (c) Despite significant differences in the context of various frames, the artifacts maintain a consistent relative position in the images (central row). Our DVT effectively denoises these artifacts, as demonstrated in the final row.
@@
With this insight, our work develops a novel two-stage denoising approach, Denoising Vision Transformers (DVT), specifically designed for removing position-dependent artifacts from pre-trained ViTs. In the first stage, we formulate a universal noise model for ViT outputs, which factorizes the output into three components: a noise-free semantics term and two terms associated with the undesirable position-based artifacts. This decomposition is achieved by enforcing cross-view feature consistency with neural fields in a per-image basis. The per-image denoising process extracts noise-free features from raw outputs and provides these clean ViT features for offline applications. In the second stage, we train a lightweight denoiser model, consisting of a single Transformer block, to predict the denoised features from the raw ViT outputs. This denoiser seamlessly integrates into pre-trained ViTs, provides denoised features for online applications, and generalizes well to unseen data.
@@
We conduct empirical evaluations to demonstrate the efficacy of DVT on seven representative ViTs: DINO [4], DINOv2 [22], DINOv2 with Register [8], DeiT-III [32], MAE [15], EVA-02 [12, 13], and CLIP [26]. These evaluations showcase significant enhancements in performance across various dense vision tasks. Our contributions are:
@@
  • We identify and highlight the widespread occurrence of noise artifacts in ViT features, pinpointing positional embeddings as a crucial underlying factor.
@@
  • We introduce a novel noise model tailored for ViT outputs, paired with a neural field-based denoising technique. This combination effectively isolates and removes noise artifacts from features.
  • We develop a streamlined and generalizable feature denoiser for real-time and robust inference.
  • Our approach significantly improves the performance of multiple pre-trained ViTs in a range of downstream tasks, confirming its utility and effectiveness (e.g., as high as a 3.84 mIoU improvement after denoising).
@@ 2 Related Works
Transformers have been used extensively across multiple domains as general-purpose feature extractors. Originally used primarily in language modeling, the Transformer architecture has found success through language-based self-training methods such as next word prediction [1, 25, 6, 33] or masked language modeling [9, 27], to name a few. In parallel, Vision Transformers pre-trained via supervised learning [35, 32, 17] or self-supervised learning [41, 15, 4, 22] have demonstrated strong generalizability to various downstream visual tasks, even without fine-tuning. In this work, we show that ViTs trained with diverse training objectives exhibit commonly observed noise artifacts in their outputs. By addressing this issue, we significantly enhance the quality of local features, as evidenced by improvements in semantic segmentation and depth prediction tasks.
@@
We study the fundamental issue of noise artifacts in ViTs, a phenomenon that has been previously noticed yet often unexplored. These artifacts are noticeable as noisy attention maps in supervised ViTs (i.e., ViTs do not attend to objects of interest well) [4, 5]. Concurrent to ours, two recent studies similarly discover artifacts even in self-supervised ViTs [39, 8]. Specifically, [8] describe these as “high-norm” patches in low-informative background regions, suggesting their occurrence is limited to large (e.g. ViT-large or greater) and sufficiently trained ViTs. However, our analysis indicates that this may not be the full picture. We find a strong correlation between the presence of artifacts and the use of positional embeddings in ViTs. This finding suggests their presence is not strictly confined to certain model sizes or training scales but is more fundamentally linked to the inherent design of ViTs. Moreover, unlike the method proposed by [8] that re-trains ViTs with register tokens [14, 38] from scratch, our approach directly denoises pre-trained models without re-training. Additionally, we note that artifacts still exist in DINOv2 trained with registers [8] (see Figure 1 DINOv2-reg, and Figure S13), and our DVT can effectively denoise them and improve their performance.
@@ 3 Preliminaries
Despite varying training approaches, the ViT architecture has largely remained consistent with its original design as presented in [10] and [35]. The forward process of a ViT, depicted in Figure 2-(b), starts by converting images into 2D patches and then embedding them, followed by a forward process of Transformer blocks. Specifically, an image $\mathbf{x}\in\mathbb{R}^{H\times W\times C}$ is first divided into patches $\mathbf{x}_{p}\in\mathbb{R}^{N\times(P^{2}\cdot C)}$, where $(H,W)$ denotes the image’s resolution, $P$ is the patch resolution, $C$ represents the number of pixel channels, and $N$ is the total number of patches. These patches are then mapped to $D$ dimensions using a trainable linear projection $\mathbf{E}\in{\mathbb{R}^{(P^{2}\cdot C)\times D}}$ to generate patch embeddings. To inject spatial information, positional embeddings, which encode patch coordinates and are denoted as $\mathbf{E}_{pos}^{i}$, are added to the patch embeddings. Formally, the forward process of a ViT is as follows:
@@
Here, $\mathbf{x}_{\text{cls}}$ and $\mathbf{E}_{pos}^{\text{cls}}$ represent the class token and its positional embedding, respectively, $L$ denotes the number of layers, and LN stands for layer normalization. Multi-head self-attention layers and multi-layer perceptron layers are termed MSA and MLP, respectively. Note that the input-independent positional embeddings operate as a spatial inductive basis and intermix with inputs, propagating through the entire ViT.
@@ 4 Denoising Vision Transformers
In this section, we start by analyzing ViT outputs to motivate our approach (§4.1). Then, we introduce our per-image denoising method, which removes artifacts and produces noise-free features (§4.2). Finally, we explain how the noise-free features are utilized as pseudo-labels to train a generalizable denoiser (§4.3). Our method pipeline is depicted in Figure 3.
Figure 3: Denoising Vision Transformers (DVT). DVT consists of a two-stage denoising pipeline. In the first stage, our method decomposes the noisy features of a crop into a noise-free semantics term $\mathcal{F}$, an input-independent, position-related artifact term $\mathcal{G}$, and an additional residual term $\Delta$ (left). In the second stage, we train a generalizable denoiser with these individually optimized, clean features (right).
@@ 4.1 Factorizing ViT Outputs
Ideal visual features should be inherently translation and reflection invariant, i.e., the features of an object should remain consistent, regardless of changes in the viewing window, size, and orientation. However, as indicated in Equations 1, 2, 3 and 4 and Figure 2-(b), ViTs intertwine patch embeddings with positional embeddings, breaking the transformation invariance of visual features. This breach of invariance might not seem immediately problematic, but our detailed investigations, as illustrated in Figure 2-(a) and (c), establish a distinct correlation between the inclusion of positional embeddings and the emergence of undesirable artifacts in ViT outputs. Particularly, the middle row of Figure 2-(c) shows that these artifacts remain nearly consistent regardless of input content, only exhibiting small residual variation across different images.
These observations motivate us to decompose ViT outputs into three terms: (1) an input-dependent, noise-free semantics term $f(\mathbf{x})$ [^1] Throughout this paper, we use “noise” and “artifacts” interchangeably.; (2) an input-independent artifact term related to spatial positions $g(\mathbf{E}_{pos})$; (3) and a residual term accounting for the co-dependency of semantics and positions $h(\mathbf{x},\mathbf{E}_{pos})$. Accordingly, we have:
$\mathrm{ViT}(\mathbf{x})=f(\mathbf{x})+g(\mathbf{E}_{pos})+h(\mathbf{x},\mathbf{E}_{pos}),$ (5)
This factorization is universally applicable to all ViTs. For instance, in scenarios where the output feature map is spatially invariant (e.g., no positional embedding is used), both $g$ and $h$ become zero functions [7]. Conversely, when every feature is dependent on both position and semantics, $f$ and $g$ turn into zero functions.
@@ 4.2 Per-image Denoising with Neural Fields
Directly addressing the above decomposition problem from a single forward pass in a ViT is impractical due to the intertwined nature of output features. To overcome this, we harness cross-view feature and artifact consistencies: (1) Feature consistency refers to the transformation invariance of visual features, wherein despite varied spatial transformations, the essential semantic content remains invariant; (2) Artifact consistency means that the input-independent artifact remains observable and constant across all transformations. Formally, consider an image $\mathbf{x}$ and a set of its randomly transformed views $T(\mathbf{x})=\{t_{0}(\mathbf{x}),t_{1}(\mathbf{x}),\cdots\}$, where each transformation $t_{i}$ is drawn from a distribution of random augmentations $\mathcal{T}$, consisting of random resizing, cropping, and flipping. Our goal is to derive a mapping $f$ that ensures the semantic features obtained from any transformed view, $f\left(t\left(\mathbf{x}\right)\right)$, remains equivalent to the transformed original semantic features, $t\left(f(\mathbf{x})\right)$. That is $f\left(t\left(\mathbf{x}\right)\right)=t\left(f(\mathbf{x})\right),t\sim\mathcal{T}$. Next, we describe our approach for jointly learning the different terms in Equation 5 to derive $f$.
Neural fields as feature mappings. At the core of our approach is to have a holistic image semantics representation, $\mathcal{F}$, for each individual image, paired with a spatial artifact feature representation, $\mathcal{G}$, shared by all transformed views. The holistic image feature representation $\mathcal{F}$ is designed to capture spatially independent, artifact-free semantics, while $\mathcal{G}$ should encode position-dependent but input-independent noise. We use neural fields [31, 20, 28, 18, 16, 39] to approximate $f$ and $g$. Specifically, we define $f(t(\mathbf{x}))=\mathcal{F}(\mathrm{coords}(t(\mathbf{x})))$, where $\mathrm{coords}(\cdot)$ extracts the pixel coordinates of the transformed views in the original image $\mathbf{x}$, and $g(\mathbf{E}^{i}_{pos})=\mathcal{G}(i)$, with $i\in\{0,\cdots,N-1\}$ denoting the patch index. For simplicity, we use $\mathcal{G}$ to denote the 2D artifact feature map reshaped from the 1D ordered sequence $\{\mathcal{G}(i)\}_{i=0}^{N-1}$. We refer to $\mathcal{F}$ and $\mathcal{G}$ as the semantics field and the artifact field, respectively.
Learning the decomposition. Our goal is to learn the semantics field $\mathcal{F}$, the artifact field $\mathcal{G}$, and the residual term $\Delta$ by minimizing a regularized reconstruction loss:
@@
Here, $\cos(\cdot,\cdot)$ denotes the cosine similarity, $\mathrm{sg}(\cdot)$ represents the stop-gradient operation, $t(\cdot)$ is a random transformation sampled from $\mathcal{T}$, and $\theta$, $\xi$ and $\psi$ are the learnable parameters. Our loss function ensures $\widehat{\Delta}$ remains minimal by imposing a sparsity regularization, thereby allowing $\widehat{\mathbf{y}^{\prime}}$ to represent as much of ViT outputs as possible. The use of stop-gradient operators is crucial to avoid trivial solutions, such as identity mapping. The reconstructed feature from our method is $\widehat{\mathbf{y}}=\mathcal{F}_{\theta}\left(\mathrm{coords}\left(t\left(\mathbf{x}\right)\right)\right)+\mathcal{G}_{\xi}+\mathrm{sg}\left(h_{\psi}\left(\mathrm{ViT}\left(t\left(\mathbf{x}\right)\right)\right)\right)$, each term corresponding to $f,g$, and $h$ as delineated in Equation 5.
Figure 4: Visual analysis of ViT output features and denoised features. (a) Visualizations of the feature maps from all layers of a DINOv2 [22] ViT-base model, using an empty image and a cat image as input. The artifacts in the cat’s feature maps have a strong visual correlation to empty input’s feature maps. (b) Visualizations of the decomposed artifacts, the original features, and the denoised features across various layers of DINOv2 ViTs. We observe similar patterns in differently-sized ViTs. (c) Visualizations of the K-Means clustering results and the cosine similarity of the central patch (red dot) to other patches. Notice that feature maps have fewer artifacts and enhanced semantic clarity after denoising, resulting in improved clustering results and similarity correspondence.
Optimization. We break our optimization process into two phases, each spanning half of the total training iterations. In the first phase, we train $\mathcal{F}_{\theta}$ and $\mathcal{G}_{\xi}$ using only $\mathcal{L}_{\text{distance}}$, allowing them to capture a significant portion of the ViT outputs. After completing half of the optimization iterations, we freeze $\mathcal{G}_{\xi}$ and continue to train $\mathcal{F}_{\theta}$ alongside $h_{\psi}$ using $\mathcal{L}_{\text{recon}}$ for the rest iterations. The coefficients $\alpha$ and $\beta$ in $\mathcal{L}_{\text{recon}}$ balance loss scales and regulate the residual term to prevent $\widehat{\Delta}$ from over-explaining the outputs.

### v2

# 5Google DeepMind equal technical contribution       project lead

@@ Abstract
We study a crucial yet often overlooked issue inherent to Vision Transformers (ViTs): feature maps of these models exhibit grid-like artifacts (“Original features” in Fig. 1), which hurt the performance of ViTs in downstream dense prediction tasks such as semantic segmentation, depth prediction, and object discovery. We trace this issue down to the positional embeddings at the input stage. To mitigate this, we propose a two-stage denoising approach, termed Denoising Vision Transformers (DVT). In the first stage, we separate the clean features from those contaminated by positional artifacts by enforcing cross-view feature consistency with neural fields on a per-image basis. This per-image optimization process extracts artifact-free features from raw ViT outputs, providing clean feature estimates for offline applications. In the second stage, we train a lightweight transformer block to predict clean features from raw ViT outputs, leveraging the derived estimates of the clean features as supervision. Our method, DVT, does not require re-training the existing pre-trained ViTs, and is immediately applicable to any Vision Transformer architecture. We evaluate our method on a variety of representative ViTs (DINO, DeiT-III, EVA02, CLIP, DINOv2, DINOv2-reg) and demonstrate that DVT consistently improves existing state-of-the-art general-purpose models in semantic and geometric tasks across multiple datasets (Fig. 1, right, Tabs. 2, 3 and 4). We hope our study will encourage a re-evaluation of ViT design, especially regarding the naive use of positional embeddings. Our code and checkpoints are publicly available in our project page.
Figure 1: Denoising Vision Transformers (DVT) effectively suppresses noisy artifacts in the visual features of all Vision Transformers (ViTs) we have tested and improves performance on a broad spectrum of dense prediction tasks, including semantic segmentation, depth estimation, object detection, and object discovery. Our evaluation encompasses a representative set of ViTs, including DINOv2 [25], DeiT-III [36], EVA-02 [13], CLIP [27], and DINOv2-reg [7]. We visualize the features before and after DVT, colored via principal component analysis (PCA). Best viewed in color. Right: We report the downstream dense prediction task performances, averaged over all models.
@@ 1 Introduction
In recent years, Transformers [38] have emerged as the universal architecture for modern foundation models across many modalities, from text [30, 6, 28, 1] to audio [20, 41], and images [9, 2]. Among these, Vision Transformers (ViTs) [9] trained at scale not only achieve state-of-the-art under multiple benchmarks but also exhibit intriguing behaviors and capabilities across various tasks [3, 16, 27, 25].
@@
Despite these significant strides made by ViTs, our work reveals a crucial yet often overlooked challenge: the presence of persistent noise artifacts in ViT outputs, observable across various training algorithms [9, 25, 36, 27, 13, 3] (illustrated in Fig. 1 left). These artifacts not only compromise visual clarity but also hinder feature interpretability and disrupt semantic coherence. For example, Fig. 2 demonstrates that applying clustering algorithms directly on the raw ViT output results in noisy clusters, and the patch feature similarity is less reliable. Additionally, these artifacts are frequently concealed by seemingly impressive performance on downstream tasks, thus evading thorough examination or detection by the research community. Addressing these artifacts can unleash the potential of pre-trained ViTs and lead to substantial performance improvements (Fig. 1 right). Therefore, our work aims to answer a crucial research question: Is it feasible to effectively denoise these artifacts in pre-trained ViTs, ideally without model retraining?
Figure 2: Artifacts hurt semantic coherence. For each triplet, we show a feature map, a K-Means cluster map, and a similarity map of the central patch (red dotted) with other patches in the image. Observe how artifacts negatively impact clustering accuracy and similarity correspondences, and how our denoising mitigates these issues.
@@
To answer this, we first investigate the origins of these artifacts. We hypothesize that positional embeddings, a fundamental component of ViT architecture, play a pivotal role in the emergence of these artifacts. Our initial analysis supports this hypothesis: First, when a zero tensor is fed into a pre-trained DINOv2 model [25], the resulting output is predominantly characterized by similar noise patterns (Fig. 3-(a, 2)). Second, we observe the absence of such artifacts in the outputs of a DINOv2 model trained without positional embeddings, which contrasts sharply with the standard model outputs (Fig. 3-(a, 1) v.s. (a, 3)). Third, take a video with continuous frames as an example (Fig. 3-(c)). Despite the significant differences in the context of various input frames, the artifacts maintain a generally consistent relative position in the images (Fig. 3-(c), middle row).
@@
With these insights, we present a two-stage denoising approach, Denoising Vision Transformers (DVT), to suppress artifacts in pre-trained ViTs. In the first stage, we obtain clean features from contaminated ones by enforcing cross-view feature consistency and artifact consistency with neural fields on a per-image basis. This per-image denoising process extracts noise-free features from raw output, providing these denoised ViT features for offline applications. In the second stage, we train a lightweight denoiser model, consisting of a single transformer block, to predict the denoised features from the raw ViT outputs. More importantly, this denoiser can be seamlessly integrated into pre-trained ViTs without extensive re-training, providing denoised features for online applications and generalizing well to unseen data.
@@
We conduct empirical evaluations to demonstrate the efficacy of DVT on six representative ViTs: DINO [3], DINOv2 [25], DINOv2 with Register (DINOv2-reg) [7], DeiT-III [36], EVA-02 [14, 13], and CLIP [27]. These evaluations demonstrate significant improvements in performance across various dense prediction vision tasks such as semantic segmentation, depth estimation, object detection, and object discovery. In summary, our contributions are:
@@
  • We identify and highlight the widespread occurrence of noise artifacts in ViT features, pinpointing positional embeddings as a crucial underlying factor. To the best of our knowledge, we are the first to provide such an analysis.
@@
  • We introduce a tailored noise model for ViTs, along with a neural field based denoising technique. This combination effectively isolates and removes noise artifacts from ViT features.
  • We develop a flexible and efficient denoiser that integrates seamlessly with pre-trained ViTs, enabling real-time applications.
  • Our approach results in substantial performance improvements across various ViTs and downstream dense prediction tasks (Fig. 1, right, Tabs. 2, 3 and 4).
Figure 3: Impact of positional embeddings in ViTs. (a) Comparison between DINOv2 ViTs [25] trained with and without positional embeddings ((“ViT” v.s. “ViT”). We show feature maps from (1) a standard ViT, (2) a ViT using only positional embeddings (PE) as input, emphasizing the emergence of artifacts, and (3) a PE-free ViT, displaying a clear absence of these artifacts. In the figure, “Patch”: patch embedding, “PE”: position embedding. (b) Illustration of how ViT retains and propagates the positional embeddings. (c) Despite significant differences in the context of various frames, the artifacts largely maintain a consistent relative position in the images (central row). Our DVT effectively denoises these artifacts, demonstrated in the final row.
@@ 2 Related Works
Transformers have been used extensively across multiple domains as general-purpose feature extractors [1, 29, 6, 37, 8, 30]. Vision Transformers [9] (ViTs) pre-trained via supervised learning [39, 36, 18] or self-supervised learning [46, 16, 3, 25] have demonstrated strong generalizability to various downstream visual tasks, even without fine-tuning. However, we show that ViTs trained with diverse training objectives exhibit commonly observed noise artifacts in their output feature maps. These artifacts are often overlooked in practice because their presence cannot be simply reflected by image classification accuracy. Thus, our work focuses on evaluating pre-trained ViTs for dense recognition tasks such as segmentation, depth estimation, and object discovery. We demonstrate how these artifacts adversely affect dense recognition tasks, thereby motivating our method to mitigate them.
@@
Our work studies the noise artifacts in ViTs, an issue that has been previously observed but often remains unexplored. These artifacts manifest as noisy attention maps in supervised ViTs (i.e., ViTs do not attend to objects of interest well) [3, 5]. Concurrently with our study, two recent studies similarly have also identified artifacts in self-supervised ViTs [44, 7]. Specifically, [7] describe these as “high-norm” patches in low-informative background regions, hypothesizing their occurrence is limited to large (e.g. ViT-large or greater) and sufficiently trained ViTs. However, our analysis indicates that this may not be the full picture, as we observe similar artifacts in small or base ViTs that cannot be easily identified by extremely high feature norm values. Instead, we find a strong correlation between the presence of artifacts and the use of positional embeddings in ViTs. This finding suggests that artifacts are not strictly confined to certain model sizes or training scales but are more fundamentally linked to the inherent design of ViTs. Moreover, unlike the method proposed by [7] that retrains ViTs with register tokens [15, 43] from scratch, our approach directly denoises pre-trained models without retraining. Users can dynamically enable or disable the plugged-in denoiser as needed. Lastly, we note that some weak artifacts still exist in DINOv2 models trained with registers [7] (see Fig. 1 DINOv2-reg and appendix), and our DVT can effectively denoise them, improving the performance of DINOv2-reg.
@@ 3 Preliminaries
Despite varying training approaches, the ViT architecture has mostly remained consistent with its original design as presented in [9] and [39]. The forward process of a ViT, depicted in Fig. 3-(b), starts by converting images into 2D patches and then embedding them, followed by a forward process of Transformer blocks. Specifically, an image $\mathbf{x}\in\mathbb{R}^{H\times W\times C}$ is first divided into patches $\mathbf{x}_{p}\in\mathbb{R}^{N\times(P^{2}\cdot C)}$, where $(H,W)$ denotes the image resolution, $P$ is the patch resolution, $C$ represents the number of pixel channels and $N$ is the number of total patches. These patches are then mapped to $D$ dimensions using a trainable linear projection $\mathbf{E}\in{\mathbb{R}^{(P^{2}\cdot C)\times D}}$ to generate patch embeddings. To inject spatial information, positional embeddings, which encode patch coordinates and are denoted $\mathbf{E}_{pos}^{i}$, are added to the patch embeddings. Formally, the forward process of a ViT is as follows:
@@
Here, $\mathbf{x}_{\text{cls}}$ and $\mathbf{E}_{pos}^{\text{cls}}$ represent the class token and its positional embedding, respectively, $L$ denotes the number of layers, and LN stands for layer normalization. Multi-head self-attention layers and multi-layer perceptron layers are termed MSA and MLP, respectively. Note how the input-independent positional embeddings function as a spatial inductive basis, intermixing with inputs and propagating throughout ViT.
@@ 4 Denoising Vision Transformers
In this section, we start by analyzing ViT outputs to motivate our approach (§4.1). Then, we introduce our per-image denoising method, which removes artifacts and produces noise-free features (§4.2). Lastly, we explain how the noise-free features are utilized as pseudo-labels to train a generalizable denoiser (§4.3). Our method pipeline is depicted in Fig. 4.
Figure 4: Method Overview. DVT consists of a two-stage denoising pipeline. (a) In the first stage, our method decomposes the raw feature of an image crop into a noise-free semantics term $\mathcal{F}$, an input-independent, position-related artifact term $\mathcal{G}$, and an additional residual term $\Delta$. (b) In the second stage, we train a generalizable denoiser to predict clean features from their original features. At inference time, only a single feedforward is needed to obtain denoised features.
@@ 4.1 Factorizing ViT Outputs
Our method is grounded in the principle that ideal visual features should be inherently translation and reflection invariant, i.e., the features of an object should remain consistent, regardless of changes in the viewing window, size, and orientation. However, as indicated in Eqs. 1, 2, 3 and 4 and Fig. 3-(b), ViTs intermix patch embeddings with positional embeddings, thereby breaking the transformation invariance of visual features. This breach of invariance might not appear immediately problematic, but our investigations, illustrated in Fig. 3-(a) and (c), reveal a distinct correlation between the inclusion of positional embeddings and the emergence of undesirable artifacts in ViT outputs. Particularly, the middle row of Fig. 3-(c) shows that these artifacts persist with minor variation across different images, highlighting their consistency independent of the input content.
These observations motivate us to decompose ViT outputs into three terms: (1) an input-dependent, noise-free semantics term $f(\mathbf{x})$ [^1] Throughout this paper, we use “noise” and “artifact” interchangeably.; (2) an input-independent artifact term related to spatial positions $g(\mathbf{E}_{pos})$; (3) and a residual term that accounts for the interdependence of semantics and positions $h(\mathbf{x},\mathbf{E}_{pos})$. The decomposition is formally expressed as:
$\mathrm{ViT}(\mathbf{x})\approx f(\mathbf{x})+g(\mathbf{E}_{pos})+h(\mathbf{x},\mathbf{E}_{pos})$ (5)
This factorization is universally applicable to all ViTs. For example, in scenarios where the output feature map is spatially invariant (e.g., no positional embedding is used), the sum of $g$ and $h$ becomes a constant bias term that can be merged into $f$.
@@ 4.2 Per-image Denoising with Neural Fields
Directly addressing the above decomposition problem within a single forward pass in a ViT is impractical due to the intertwined nature of output features. To overcome this, we exploit the consistencies in cross-view features and artifacts: (1) Feature consistency refers to the transformation invariance of visual features, where despite the varied spatial transformations, the semantic content remains invariant; (2) Artifact consistency means that the input-independent artifact remains observable and constant across all transformations. Formally, consider an image $\mathbf{x}$ and a set of its randomly transformed views $T(\mathbf{x})=\{t_{0}(\mathbf{x}),t_{1}(\mathbf{x}),\cdots\}$, where each transformation $t_{i}$ is drawn from a distribution of random augmentations $\mathcal{T}$, consisting of random resizing, cropping, and flipping. Our goal is to derive a mapping $f$ such that the semantic features obtained from any transformed view, $f\left(t\left(\mathbf{x}\right)\right)$, are equivalent to the transformed original semantic features, $t\left(f(\mathbf{x})\right)$; that is, $f\left(t\left(\mathbf{x}\right)\right)=t\left(f(\mathbf{x})\right)$ with $t\sim\mathcal{T}$. Next, we describe our approach for learning the different terms in Eq. 5 in conjunction to derive $f$.
Neural fields as feature mappings.
At the core of our approach is to have a holistic image semantics representation $\mathcal{F}$, for each individual image, alongside a spatial artifact feature representation, $\mathcal{G}$, shared by all transformed views. The holistic image feature representation $\mathcal{F}$ is designed to capture spatially independent, artifact-free semantics, while $\mathcal{G}$ should encode position-dependent but input-independent noise. We use coordinate networks, known as neural fields [35, 23, 32, 19, 17, 44], to actualize $\mathcal{F}$ and $\mathcal{G}$. Specifically, we define $f(t(\mathbf{x}))=\mathcal{F}(\mathrm{coords}(t(\mathbf{x})))$, where $\mathrm{coords}(\cdot)$ extracts the pixel coordinates of the transformed views relative to the original image $\mathbf{x}$, and $g(\mathbf{E}^{i}_{pos})=\mathcal{G}(i)$, with $i\in\{0,\cdots,N-1\}$ denoting the patch index. For simplicity, we use $\mathcal{G}$ to denote the 2D artifact feature map reshaped from the 1D ordered sequence $\{\mathcal{G}(i)\}_{i=0}^{N-1}$. We refer to $\mathcal{F}$ and $\mathcal{G}$ as the semantics field and the artifact field, respectively.
Learning the decomposition.
We learn the semantics field $\mathcal{F}$, the artifact field $\mathcal{G}$, and the residual term $\Delta$ by minimizing a regularized reconstruction loss:
@@
Here, $\cos(\cdot,\cdot)$ denotes the cosine similarity, $\mathrm{sg}(\cdot)$ represents the stop-gradient operation, $t(\cdot)$ is a random transformation sampled from $\mathcal{T}$, and $\theta$, $\xi$ and $\psi$ are the learnable parameters. Our loss function is designed to encourage $\widehat{\Delta}$ to remain minimal by imposing a sparsity regularization, thereby allowing $\widehat{\mathbf{y}^{\prime}}$ to represent as much of the ViT output as possible. The use of stop-gradient operators is to avoid trivial solutions, such as identity mapping. The reconstructed feature from our method is $\widehat{\mathbf{y}}=\mathcal{F}_{\theta}\left(\mathrm{coords}\left(t\left(\mathbf{x}\right)\right)\right)+\mathcal{G}_{\xi}+\mathrm{sg}\left(h_{\psi}\left(\mathrm{ViT}\left(t\left(\mathbf{x}\right)\right)\right)\right)$, each term corresponding to $f,g$, and $h$ as defined in Eq. 5.
Optimization.
We break our optimization process into two phases, each spanning half of the total training iterations. In the first phase, we train $\mathcal{F}_{\theta}$ and $\mathcal{G}_{\xi}$ using only $\mathcal{L}_{\text{distance}}$, allowing them to capture a significant portion of the ViT outputs. After completing half of the optimization iterations, we freeze $\mathcal{G}_{\xi}$ and continue to train $\mathcal{F}_{\theta}$ alongside $h_{\psi}$ using $\mathcal{L}_{\text{recon}}$ for the rest iterations. The coefficients $\alpha$ and $\beta$ in $\mathcal{L}_{\text{recon}}$ balance loss scales and regulate the residual term to prevent $\widehat{\Delta}$ from over-explaining the outputs.